OpenStudy (anonymous):

Li has to read four books from a list of 10 books. How many sets of four books can be made? 24 210 7

OpenStudy (kinggeorge):

This is a combinations question where you're asked to find the value of \(10C4\) or\[\binom{10}{4} ={10! \over 4!(10-4)!}\]

OpenStudy (anonymous):

Thanks i forgot the - part

OpenStudy (kinggeorge):

That would help :P

OpenStudy (anonymous):

umm i got 25200 i still need help

OpenStudy (zarkon):

that is a little large :)

OpenStudy (zarkon):

from what KG gave you \[\binom{10}{4} ={10! \over 4!(10-4)!}=\frac{10*9*8*7}{4*3*2*1}=210\]

OpenStudy (anonymous):

yeah when i canceled out the 4-3-2-1 on both and subtracted 10 by 4 the divided 151200 by 6 i got 25200

OpenStudy (anonymous):

ohh opps thanks alot i own ya one